# Junzhou Huang8, Faez Ahmed9, Akash Srivastava2 & Dimitris N. Metaxas1 1

### v1

@@ 4.1 Image diffusion model
For the image diffusion model, we mainly investigate the use of absorbing state discrete model (Austin et al., 2021) including a masked generative model, Paella, and a multinomial diffusion model, VQ-Diffusion. We demonstrate the inversion reconstruction ability and image editing performance in both categories with DICE.
Dataset. The Prompt-based Image Editing Benchmark (PIE-Bench) by  (Ju et al., 2023) is a recently introduced dataset designed to evaluate text-to-image (T2I) editing methods. The dataset assesses language-guided image editing in 9 different scenarios with 700 images. The benchmark’s detailed annotations and variety of editing tasks were instrumental in thoroughly assessing our method’s capabilities, ensuring a fair and consistent comparison with existing approaches.
@@ 4.1.1 Inversion Reconstruction
Evaluation Metrics. Here, we evaluate the image similarity by PSNR, LPIPS, MSE and SSIM of the original and the generated image under the same prompt with DICE and masked generation.
@@
Method Metric
Inverse+Model PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
Inpainting+Paella 10.50 565.11 1002.09 30.13
Ours+Paella 30.91 39.81 11.07 90.22
$\text{Ours}^{\dagger}$+Paella Inf 0.07 0.01 99.99
Table 1: Inversion Reconstruction performance $\dagger$ The metric is calculated between the original image and its inverted counterpart. Due to the encoding and decoding steps in the VQ-VAE/GAN process, some inaccuracies are introduced by the quantization. The PSNR is Inf due to the reconstruction of our method yielding the same image after the VQ-VAE/GAN process.
4.1.2 Editing Performance
Evaluation Metrics. To demonstrate the effectiveness and efficiency of our proposed inversion method, we employ eight metrics covering three key aspects: structure distance, background preservation, and edit prompt-image consistency, as outlined in Ju et al. (2023). We utilize the structure distance metric proposed by Tumanyan et al. (2023) to measure the structural similarity between the original and generated images. To evaluate how well the background is preserved outside the annotated editing mask, we use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Mean Squared Error (MSE), and Structural Similarity Index Measure (SSIM) (Wang et al., 2004). We also assess the consistency between the edit prompt and the generated image using CLIP (Radford et al., 2021) Similarity Score  (Wu et al., 2021), which is calculated over the whole image and specifically within the regions defined by the editing mask.
Figure 4: Visualization of editing results. Editing results for our method using Paella and VQ-Diffusion are presented, along with their corresponding prompts. The results demonstrate that our method can effectively modify the input image according to the target prompt while preserving the image structure. Editing with masked generative model (Paella (Rampas et al., 2022)) is more stable and easier than with multinomial diffusion models (VQ-Diffusion (Gu et al., 2022)).
@@ 4.1.2 Editing Performance
Table 2: Quantitative results on image editing performance. Comparison of our proposed method with the masked inpainting with the discrete diffusion model Paella, as well as continuous diffusion model (Stable Diffusion v1.4) using DDIM inversion. “P2P” refers to Prompt-to-Prompt (Hertz et al., 2022), and “Prompt” denotes editing performed solely through forward edit prompts. Entries marked with an asterisk () are cited from Ju et al. (2023). : For VQ-Diffusion, the images are down-sampled to $256\times 256$. It is important to note that due to differences in base models and editing algorithms, the metrics across methods are not directly comparable. However, our method significantly outperforms both inpainting and strong baselines (e.g., Null-Text Inversion + SD1.4) in terms of structural preservation. As expected, inpainting achieves a high CLIP score since it directly generates image patches based on the target prompt.
@@
Inverse Editing PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
@@ 4.2 Language Diffusion Model
In this section, we evaluate DICE on RoBERTa (Liu et al., 2019), a text discrete diffusion model, to generate sentences with opposing sentiments while preserving structural similarities. We begin with two prompts—one with a positive sentiment and another with a negative sentiment. Each prompt contains two sentences: the first sentence indicates the sentiment type and sets the contextual background, and the second sentence is the target for inversion and generation. Initially, we invert the second sentence of the negative sentiment prompt using the entire prompt as context, which produces a noised token representation of that sentence. Next, we condition the model on the positive sentiment by concatenating the first sentence of the positive sentiment prompt with the noised token of the inverted negative sentence. This setup guides the model to generate a new second sentence that mirrors the structure of the original negative sentence but expresses a positive sentiment instead. Through this process, we assess the model’s capability to invert and generate text that aligns with a specified sentiment while retaining the original sentence’s structural elements.
@@
Inversion Process. In our experiment, we specifically focus on inverting the second sentence, indicated as red in Table 6, while keeping the first sentence intact (black), as it usually contains essential context. During the reverse process, we aim to reconstruct/edit the second sentence by recovering it from the noised tokens acquired in the inversion phase.
Dataset 1. Positive Sentiment: Thanks to her efforts. The event was a huge success. Negative Sentiment: Despite her efforts. The event was a complete disaster. 2. Positive Sentiment: This book is definitely interesting. I can’t put it down; it’s full of surprises. Negative Sentiment: This book is definitely interesting. I can’t wait to finish it; it’s so predictable. 3.
@@
Dataset Generation. In order to evaluate the editing performance, we designed and proposed a new dataset called Sentiment Editing. The objective is to edit the sentiment of the sentence while preserving the structure of the sentence and also sticking to the theme of the sentence. Here, we demonstrate two sets of sentences in our dataset. Please refer to supplementary materials for the process of generating the dataset and more examples.
4.2.1 Inversion Reconstruction
Evaluation Metric. For reconstruction, we use Hit Rate, which is defined as the proportion of cases where each method generates an identical sentence to the original. In addition, we compute the Semantic Textual Similarity (STS) score by measuring the cosine similarity between the sentence embeddings, using the model proposed by Reimers (2019) et al.
Method Metric
Inverse+Model Accuracy×${}_{{}^{\times 10^{2}}}$uparrow Textual Similarity×${}_{{}^{\times 10^{2}}}$uparrow
Masked Generation+RoBERTa 0.0 6.57
Ours+RoBERTa 99.74 99.90
Method Metric
Inverse+Model Structure Preservation×${}_{{}^{\times 10^{2}}}$uparrow Sentiment Correctness×${}_{{}^{\times 10^{2}}}$uparrow
Masked Generation+RoBERTa 29.80 12.94
Ours+RoBERTa 94.76 72.51
4.2.2 Sentence Editing
In this section, we evaluate the editing performance of the proposed inversion method on RoBERTa. In Table 6, the sentence shown in black under the negative prompt column is input during the inversion process. The sentence that is being inverted is displayed in red. For editing, the prompt is then substituted with the black sentence on the right, and noise is added at the end for the forward process. The output of the forward process for the noise is presented in blue.
Evaluation Metric. For the sentence editing task, we evaluate the generated sentences based on two criteria: (1) structural preservation, which assesses whether the sentence structure is retained, and (2) sentiment correctness, which evaluates whether the sentiment of the edited sentence aligns with the sentiment of the original prompt. Both the structural preservation rate and sentiment correctness rate are calculated using ChatGPT-4 (Achiam et al., 2023) as a classifier. The details of using ChatGPT for evaluation can be reviewed in Supplementary Materials.
Results. Table 5 presents a comparative analysis of two text editing methods that both employ RoBERTa, focusing on the effectiveness in terms of Structure Preservation and Sentiment Correctness. Our method significantly outperforms masked generation in both metrics. This difference highlights the superior capability of our inversion method to encode the original structure of the text in the latent space and the flexibility to adjust its sentiment more accurately. In Table 6, we demonstrate both the initial prompt and the edited result. Our approach retains the sentence structure of the negative prompt while modifying its sentiment to a more positive one.
@@ 4.2.2 Sentence Editing
Negative Sentiment: The new office space is fantastic. Positive Sentiment: The new office space is fantastic.
@@ 5 Conclusion
In this paper, we introduced DICE (Discrete Inversion for Controllable Editing), an inversion algorithm for discrete diffusion models, including multinomial diffusion and masked generative models. By leveraging recorded noise sequences and masking patterns during the reverse diffusion process, DICE enables accurate reconstruction and flexible editing of discrete data without the need for predefined masks or cross-attention manipulation. Our experiments across multiple models and modalities, such as images and text, demonstrate the effectiveness of DICE in preserving data fidelity while enhancing editing capabilities. Furthermore, we demonstrate the potential of DICE for converting RoBERTa, a model traditionally focused on data understanding, into a generative model for text generation and editing. We believe that DICE enhances the capabilities of discrete generative models, offering new opportunities for fine-grained content manipulation in discrete spaces.
Appendix A Details on Multinomial Diffusion Models
@@
Appendix B Analysis on Mutual Information
Proof of Remark 3.1.
@@ Appendix B Analysis on Mutual Information
$=(\sqrt{\overline{\alpha}_{t-1}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t-1}}\bm{\epsilon})-\sqrt{\alpha_{t}}(\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}^{\prime})$
@@
$=\frac{D}{2}\log(2\pi e(\beta_{t}^{2}\overline{\alpha}_{t-1}+1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t}))-\frac{D}{2}\log(2\pi e(1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t}))$
@@
Appendix C Implementation Details
For all reconstruction task, we employ a $\tau=1.0$ and $\lambda_{1}=1.0,\lambda_{2}=0.0$ with 32 sampling steps and 26 renoising steps. For editing tasks, the hyper-parameters are summarized in Table 7.
Editing Experiment Hyper-parameters
Method Configuration CFG $\lambda_{1}$ $\lambda_{2}$ $\tau$
Paella Set 1 10.0 0.7 0.3 0.9
VQ-Diffusion Set 1 5.0 0.2 0.8 1.0
RoBERTa Sentiment Set 1 - 0.2 0.8 0.7
Set 2 - 0.25 0.75 0.7
Table 7: Hyper-parameters for Paella, VQ-Diffusion, and RoBERTa sentiment editing experiments. For sentiment editing task with RoBERTa, we utilize two sets of hyper-parameters empirically due to the variance in the sentence length.
@@ Appendix C Implementation Details
Appendix D Ablation Studies
D.1 Noise Injection Function
@@
$\operatorname*{arg\,max}{(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}+\lambda_{2}\cdot\bm{g})}=\operatorname*{arg\,max}{(\frac{1}{\lambda_{2}}\left(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}\right)+\bm{g})},~{}~{}\lambda_{2}>0.$
@@
Max. The third way is inspired by the property of Gumbel distribution (Wikipedia contributors, 2024), that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@ D.2 Hyperparameter Search
Appendix E Additional Results on Image Editing
@@
Appendix F Additional Results on Text Editing
@@
  1. 1.
    Positive Sentiment: Thanks to her efforts. The event was a huge success.
    Negative Sentiment: Despite her efforts. The event was a complete disaster.
  2. 2.
    Positive Sentiment: This book is definitely interesting. I can’t put it down; it’s full of surprises.
    Negative Sentiment: This book is definitely interesting. I can’t wait to finish it; it’s so predictable.
  3. 3.
    Positive Sentiment: The new office space is fantastic. It’s spacious and perfect for productivity.
    Negative Sentiment: The new office space is fantastic. It’s cramped and lacks proper facilities.
  4. 4.
    Positive Sentiment: Thanks to her efforts. The event was a huge success.
    Negative Sentiment: Despite her efforts. The event was a complete disaster.
  5. 5.
    Positive Sentiment: Regarding the lecture. It was insightful and engaging.
    Negative Sentiment: Regarding the lecture. It was dull and confusing.
  6. 6.
    Positive Sentiment: Despite the initial problems. The project was a success.
    Negative Sentiment: Despite the initial problems. The project ended in failure.
  7. 7.
    Positive Sentiment: Regarding the new app. It’s user-friendly and very helpful.
    Negative Sentiment: Regarding the new app. It’s complicated and not useful.
  8. 8.
    Positive Sentiment: Reflecting on my environmental initiatives. Implementing changes has reduced my carbon footprint.
    Negative Sentiment: Reflecting on my environmental initiatives. It’s challenging to maintain, and progress is slow.
  9. 9.
    Positive Sentiment: The business proposal was well-received. The ideas were innovative, and the presentation was convincing.
    Negative Sentiment: The business proposal was rejected. The ideas were impractical, and the presentation was unconvincing.
  10. 10.
    Positive Sentiment: The training program was highly effective. It boosted skills and confidence, and everyone left motivated.
    Negative Sentiment: The training program was ineffective. It didn’t teach much, and most people left feeling unmotivated.
  11. 11.

### v2

@@ 4.1 Image diffusion model
For the image diffusion model, we mainly investigate the use of absorbing state discrete model [2] including a masked generative model, Paella, and a multinomial diffusion model, VQ-Diffusion. We demonstrate the inversion reconstruction ability and image editing performance in both categories with DICE.
Dataset. The Prompt-based Image Editing Benchmark (PIE-Bench) by  [28] is a recently introduced dataset designed to evaluate text-to-image (T2I) editing methods. The dataset assesses language-guided image editing in 9 different scenarios with 700 images. The benchmark’s detailed annotations and variety of editing tasks were instrumental in thoroughly assessing our method’s capabilities, ensuring a fair and consistent comparison with existing approaches.
Figure 3: Visualization of editing results. Editing results for our method using Paella and VQ-Diffusion are presented, along with their corresponding prompts. The results demonstrate that our method can effectively modify the input image according to the target prompt while preserving the image structure. Editing with masked generative model (Paella [43]) is more stable and easier than with multinomial diffusion models (VQ-Diffusion [16]).
@@ 4.1.1 Inversion Reconstruction
Evaluation Metrics. Here, we evaluate the image similarity by PSNR, LPIPS, MSE and SSIM of the original and the generated image under the same prompt with DICE.
@@
Table 2: Quantitative results on image editing performance. Comparison of our proposed method with the masked inpainting with the discrete diffusion model Paella, as well as continuous diffusion model (Stable Diffusion v1.4) using DDIM inversion. “P2P” refers to Prompt-to-Prompt [21], and “Prompt” denotes editing performed solely through forward edit prompts. Entries marked with an asterisk () are cited from [28]. : For VQ-Diffusion, the images are down-sampled to $256\times 256$. It is important to note that due to differences in base models and editing algorithms, the metrics across methods are not directly comparable. However, our method significantly outperforms both inpainting and strong baselines (e.g., Null-Text Inversion + SD1.4) in terms of structural preservation. As expected, inpainting achieves a high CLIP score since it directly generates image patches based on the target prompt.
4.1.2 Editing Performance
Evaluation Metrics. To demonstrate the effectiveness and efficiency of our proposed inversion method, we employ eight metrics covering three key aspects: structure distance, background preservation, and edit prompt-image consistency, as outlined in [28]. We utilize the structure distance metric proposed by [51] to measure the structural similarity between the original and generated images. To evaluate how well the background is preserved outside the annotated editing mask, we use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [58], Mean Squared Error (MSE), and Structural Similarity Index Measure (SSIM) [53]. We also assess the consistency between the edit prompt and the generated image using CLIP [41] Similarity Score  [55], which is calculated over the whole image and specifically within the regions defined by the editing mask.
@@ 4.1.2 Editing Performance
Inversion+Model Editing PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
@@ 4.2 Language Diffusion Model
In this section, we evaluate DICE on RoBERTa [31], a text discrete diffusion model, to generate sentences with opposing sentiments while preserving structural similarities. We begin with two prompts—one with a positive sentiment and another with a negative sentiment. Each prompt contains two sentences: the first sentence indicates the sentiment type and sets the contextual background, and the second sentence is the target for inversion and generation. Initially, we invert the second sentence of the negative sentiment prompt using the entire prompt as context, which produces a noised token representation of that sentence. Next, we condition the model on the positive sentiment by concatenating the first sentence of the positive sentiment prompt with the noised token of the inverted negative sentence. This setup guides the model to generate a new second sentence that mirrors the structure of the original negative sentence but expresses a positive sentiment instead. Through this process, we assess the model’s capability to invert and generate text that aligns with a specified sentiment while retaining the original sentence’s structural elements.
@@
Inversion Process. In our experiment, we focus on inverting the second sentence, indicated as red in the dataset, while keeping the first sentence intact (black), as it usually contains essential context. During the reverse process, we aim to reconstruct/edit the second sentence by recovering it from the noised tokens acquired in the inversion phase.
@@
Dataset Generation. In order to evaluate the editing performance, we designed and proposed a new dataset for Sentiment Editing. The objective is to edit the sentiment of the sentence while preserving the structure of the sentence and also sticking to the theme of the sentence. Here, we demonstrate two sets of sentences in our dataset. Please refer to supplementary materials for the details.
@@
Table 4: Editing results of our method with RoBERTa. The sentence in red is the one being inverted, and the blue sentence represents the editing result.
4.2.1 Inversion Reconstruction
Evaluation Metric. For reconstruction, we use Hit Rate, which is defined as the proportion of cases where each method generates an identical sentence to the original. In addition, we compute the Semantic Textual Similarity (STS) score by measuring the cosine similarity between the sentence embeddings, using the model proposed by [44].
Editing Method Accuracy×${}_{{}^{\times 10^{2}}}$uparrow Textual Similarity×${}_{{}^{\times 10^{2}}}$uparrow
Masked Generation 0.0 6.57
Ours 99.74 99.90
Editing Method Structure Preservation×${}_{{}^{\times 10^{2}}}$uparrow Sentiment Correctness×${}_{{}^{\times 10^{2}}}$uparrow
Masked Generation 29.80 12.94
Ours 94.76 72.51
4.2.2 Sentence Editing
In this section, we evaluate the editing performance of the proposed inversion method on RoBERTa.
Evaluation Metric. For the sentence editing task, we evaluate the generated sentences based on two criteria: (1) structural preservation, which assesses whether the sentence structure is retained, and (2) sentiment correctness, which evaluates whether the sentiment of the edited sentence aligns with the sentiment of the original prompt. Both the structural preservation rate and sentiment correctness rate are calculated using ChatGPT-4 [1] as a classifier. Qualitative samples are given in Table 7.
Results. Table 6 presents a comparative analysis of two text editing methods that both employ RoBERTa, focusing on the effectiveness in terms of Structure Preservation and Sentiment Correctness. Our method significantly outperforms masked generation in both metrics. This difference highlights the superior capability of our inversion method to encode the original structure of the text in the latent space and the flexibility to adjust its sentiment more accurately. In Supplementary Materials, we demonstrate both the initial prompt and the edited result. Our approach retains the sentence structure of the negative prompt while modifying its sentiment to a more positive one.
@@ 5 Conclusion
In this paper, we introduced DICE, an inversion algorithm for discrete diffusion models, including multinomial diffusion and masked generative models. By leveraging recorded noise sequences and masking patterns during the reverse diffusion process, DICE enables accurate reconstruction and flexible editing of discrete data without the need for predefined masks or cross-attention manipulation. Our experiments across multiple models and modalities demonstrate the effectiveness of DICE in preserving data fidelity while enhancing editing capabilities. Furthermore, we demonstrate the potential of DICE for converting RoBERTa, a model traditionally focused on data understanding, into a generative model for text generation and editing. We believe that DICE enhances the capabilities of discrete generative models, offering new opportunities for fine-grained content manipulation in discrete spaces.
Acknowledgment. This research has been partially funded by research grants to Dimitris N. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, FA9550-23-1-0417, and NIH-5R01HL127661.
Appendix
Figure 4: This paper was accepted to CVPR 2025 but later desk-rejected post camera-ready.
Appendix A
Figure 4: This paper was accepted to CVPR 2025 but later desk-rejected post camera-ready, due to a withdrawal from ICLR made 14 days before reviewer assignment.
Appendix B Details on Multinomial Diffusion Models
@@
Appendix C Analysis on Mutual Information
Here we provide an analysis to quantify the amount of information encoded in latent. Since the inversion involves model forward function call which is difficult to analyze. We describe in the following a simple yet prototypical example of DDPM, where the posterior mean can be computed in closed-form thus allows us to compute the mutual information.
Remark C.1.
Given a simple Gaussian DDPM with ${\bm{x}}_{0}\sim\mathcal{N}({\bm{0}},{\bm{I}})$, latents $\{{\bm{z}}_{t}\}$ are obtained with DDPM inversion [26], then the mutual information between ${\bm{z}}_{t}$ and ${\bm{x}}_{0}$:
The mutual information between $\bm{z}_{t}$ and $\bm{x}_{0}$ is illustrated in Supplementary Materials. We observe that the amount of information encoded from $\bm{x}_{0}$ into $\bm{z}_{t}$ decreases as $t$ increases, motivating us to explore different scheduling strategies for $\lambda$’s.
@@ Appendix C Analysis on Mutual Information
$=(\sqrt{\overline{\alpha}_{t-1}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t-1}}\bm{\epsilon})$
$-\sqrt{\alpha_{t}}(\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}^{\prime})$
@@
$=\frac{D}{2}\log(2\pi e(\beta_{t}^{2}\overline{\alpha}_{t-1}+1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t}))$
$-\frac{D}{2}\log(2\pi e(1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t}))$
@@
We also provide the relationship between the mutual information of $z_{t},z_{0}$ and the timestep $t$ in Fig. 5.
Appendix D Implementation Details
For all reconstruction task, we employ a $\tau=1.0$ and $\lambda_{1}=1.0,\lambda_{2}=0.0$ with 32 sampling steps and 26 renoising steps.
The hyper-parameters for Paella editing experiment is CFG$=10.0$, $\lambda_{1}=0.7$, $\lambda_{2}=0.3$ and $\tau=0.9$. The hyper-parameters for VQ-Diffusion in editing is CFG$=5.0$, $\lambda_{1}=0.2$, $\lambda_{2}=0.8$.
For sentiment editing task with RoBERTa, we utilize two sets of hyperparameter: $\tau=0.7$, $\lambda_{1}=0.2$, $\lambda_{2}=0.8$ and $\tau=0.7$, $\lambda_{1}=0.25$, $\lambda_{2}=0.75$.
@@ Appendix D Implementation Details
Appendix E Ablation Studies
E.1 Noise Injection Function
@@
$\operatorname*{arg\,max}{(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}+\lambda_{2}\cdot\bm{g})}$
$=$ $\operatorname*{arg\,max}{(\frac{1}{\lambda_{2}}\left(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}\right)+\bm{g})},~{}~{}\lambda_{2}>0.$
@@
Max. The third way is inspired by the property of Gumbel distribution [54], that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@ E.2 Hyperparameter Search
Appendix F Additional Results on Image Editing
@@
Additional baselines. We compare with SDEdit [35] and ControlNet [57] [^1] We use the ControlNet-InPaint model based on Stable Diffusion v1.5: https://github.com/mikonvergence/ControlNetInpaint. Results are shown in Figure 13 and Table 8.
Noise injection functions. We compare various noise injection functions, including taking the maximum of Gumbel noise and the recorded noise, as well as the variance-preserving noise injection function.
Mask schedule functions. In Figure 14, we present four types of mask scheduling functions: (a, c) concave up and (b, d) concave down. Our results indicate that concave up mask scheduling functions perform better than their concave down counterparts. Quantitative results are shown in Table 9.
Comparison between inclusive and random masks. To understand the impact of randomness in the masking schedule, we illustrate masks that are inclusive compared to totally random. Inclusive mask is mask schedule that are increasingly growing, which is used in Paella, compared to randomly sampled masks.
Appendix G Latent Space Properties
Figure 9: Latent space properties of DICE.
Figure 9 demonstrates additional experiment results on the latent space property including flipping and shifting in the latent space.
Appendix H Details on Text Editing Experiments
@@
Dataset 1. Positive Sentiment: Thanks to her efforts. The event was a huge success. Negative Sentiment: Despite her efforts. The event was a complete disaster. 2. Positive Sentiment: This book is definitely interesting. I can’t put it down; it’s full of surprises. Negative Sentiment: This book is definitely interesting. I can’t wait to finish it; it’s so predictable. 3. Positive Sentiment: Regarding the lecture. It was insightful and engaging. Negative Sentiment: Regarding the lecture. It was dull and confusing. 4. Positive Sentiment: Despite the initial problems. The project was a success. Negative Sentiment: Despite the initial problems. The project ended in failure. 5. Positive Sentiment: Regarding the new app. It’s user-friendly and very helpful. Negative Sentiment: Regarding the new app. It’s complicated and not useful. 6. Positive Sentiment: Reflecting on my environmental initiatives. Implementing changes has reduced my carbon footprint. Negative Sentiment: Reflecting on my environmental initiatives. It’s challenging to maintain, and progress is slow. 7. Positive Sentiment: The business proposal was well-received. The ideas were innovative, and the presentation was convincing. Negative Sentiment: The business proposal was rejected. The ideas were impractical, and the presentation was unconvincing. 8. Positive Sentiment: The training program was highly effective. It boosted skills and confidence, and everyone left motivated. Negative Sentiment: The training program was ineffective. It didn’t teach much, and most people left feeling unmotivated. 9.
Appendix I Additional Results on Sentiment Editing
Negative Prompt Our Edited Results
It was dull and confusing. It was clear and surprising.
The project ended in failure. New project still in progress.
It’s complicated and not useful. It’s On and It’s Epic.